OpenStudy (anonymous):

limit ( {(a+1)^n - a^n) where n tends to inifnity assuming that a>0 but also consider a<0 and a>=1.

OpenStudy (dls):

\[\LARGE \lim_{n \rightarrow \infty}(a+1)^{n}-a^n \] you mean this?

OpenStudy (anonymous):

The two terms must be included in braces meaning that the limit of the difference of the two must be determined.

OpenStudy (dls):

\[\LARGE \lim_{n \rightarrow \infty}((a+1)^{n}-a^n)\]

OpenStudy (anonymous):

how about using this identity\[b^n-a^n=(b-a)(b^{n-1}+b^{n-2}a+...ba^{n-2}+a^{n-1})\]are u familiar with that?

OpenStudy (anonymous):

No not.Please give more information

OpenStudy (anonymous):

\(\begin{align*}(a+1)^n&=a^n+a^{n-1}+a^{n-2}+\cdots+a^2+a+1\\ (a+1)^n-a^n&=a^{n-1}+a^{n-2}+\cdots+a^2+a+1\end{align*}\) Note that the RHS can be written as \[\large\sum_{k=0}^{n-1}a^k\] Since you're taking the limit as n approaches infinity, you basically have \[\large\sum_{k=0}^{\infty}a^k\] Now let's use what we know about a geometric series like this one. For \(a\ge1,\) the series diverges. For \(0<a<1,\) the series converges. For \(a=0,\) the terms are all 0, so the sum is 0 (convergent). For \(-1<a<0,\) the series converges. For \(a\le-1,\) the series diverges.

OpenStudy (anonymous):

Thank you very much. This is well explained.

OpenStudy (anonymous):

oh oh, be careful\[(1+a)^n\neq a^n+a^{n-1}+...+a+1\]

OpenStudy (anonymous):

@Hennie, make note of what @mukushla pointed out. I forgot the coefficients on most of the terms, but you can still write it as a series. Thanks for the correction